What is the equation of the line going through the points (5,6) and (10, 9).
Answer by mananth(16949)

x1 y1 x2 y2
5 6 10 9

slope m = (y2-y1)/(x2-x1)
(9-6)/(10-5)
(3/5)
m=-3

Plug value of the slope and point (7,9) in
Y = m x + b
9.00 = -21 + b
b= 9 -21
b= 30
So the equation will be
Y = -3 x + 30
m.ananth@hotmail.ca